Prove or disapprove that the complex sphere is compact.


1
Expert's answer
2021-08-16T16:57:02-0400

The complex sphere "\bar\mathbb{C}" is a complex manifold with two charts "(U_1,z_1)" and "(U_2,z_2)", both with domain equal to the complex number plane "\mathbb{C}". The intersection of domains "U_1" and "U_2" is

"U_1\cap U_2=U_1\setminus\{z_1=0\}=U_2\setminus\{z_2=0\}", the transition map "z_2=z_1^{-1}".

Let "K_1=\{z_1\in U_1:|z_1|\leq 1\}", "K_2=\{z_2\in U_2:|z_2|\leq 1\}" be two closed unit disks. They are compact as closed bounded subsets of "\mathbb{C}".

We will prove that the complex sphere "\bar\mathbb{C}" is compact by showing that "\bar\mathbb{C}=K_1\cup K_2". It is evident that "\bar\mathbb{C}\supset K_1\cup K_2".

Consider any point "P\in U_1\cap U_2".

If "z_1(P)\leq 1", then "P\in K_1".

If "|z_1(P)|> 1", then "|z_2(P)|=|z_1(P)|^{-1}\leq 1" and "P\in K_2". Therefore, "U_1\cap U_2\subset K_1\cup K_2".

Since the point "\{z_1=0\}\in K_1" and the point "\{z_2=0\}\in K_2", it follows that"\bar\mathbb{C}=U_1\cup U_2=(U_1\cap U_2)\cup\{z_1=0\}\cup\{z_2=0\}\subset K_1\cup K_2".

Therefore "\bar\mathbb{C}=K_1\cup K_2" and "\bar\mathbb{C}" is compact.